kekeman:

How many solutions exist for the equation cos 2θ − sin2θ = 0 on the interval [0, 360°)? 5 4 3 2

kekeman:

I need some help

surjithayer:

\[\cos 2\theta-\sin2 \theta=0,\cos 2\theta=\sin 2\theta\] \[\tan 2\theta=1=\tan 45,\tan (180+45)\] \[=\tan45,\tan 225\] \[=\tan \left( 45+180n \right),\tan (225+180n),n=0,1\] \[2\theta=45+180n,225+180n,n=0,1\] when n=0 \[2\theta=45+180\times 0,225+180\times 0\] \[2\theta=45,225\] \[\theta=\frac{ 45 }{ 2 },\frac{ 225 }{ 2 }=22.5,112.5\] when n=1 \[2\theta=45+180\times1,225+180\times 1\] \[2\theta=225,405\] \[\theta=112.5,202.5\] when n=2 \[2\theta=45+180\times2,225+180\times 2\] \[2\theta=45+360,225+360\] \[2\theta =405,585\] \[\theta=202.5,292.5\] when n=3 \[2\theta=45+180\times3,225+180\times 3\] \[2\theta=45+540,225+540\] \[2\theta=585,765\] \[\theta=292.5,382.5\] \[\theta=22.5,112.5,202.5,292.5\] so there are four solutions.

kekeman:

Okay wow this makes so much sense great job, thank you so much!

surjithayer:

n=0,1,2,3